OpenStudy (anonymous):

Find the quadratic equation for roots 1/3 and -1/4

OpenStudy (anonymous):

Here's the process. If your roots are... say, 1 and 2, then the equations for these roots are x=1 and x=2. In order to get a quadratic equation, all numbers and variables must be on the same side of the equation. So in both cases, subtract the number over to the left. x-1=0 and x-2=0 Both of these are now your zeroes. These now become your factors for your quadratic equation. (x-1)(x-2)=0 Then you multiply them out using the FOIL method: Multiply the front numbers together, the outside numbers together, the inside numbers together, and the last numbers together. x times x, x times -2, -1 times x, -1 times -2. then you add them together. x^2+(-2x)+(-x)+2=0 This comes out to be x^2-3x+2=0 And that is the equation.

OpenStudy (amoodarya):

quadratic equation for roots x1,x2 is (x-x1)*(x-x2)=0 like this quadratic equation for roots 2,4 (x-2)(x-4)=x^2-6x+8

OpenStudy (whpalmer4):

Any polynomial with rational coefficients and roots \(r_1,r_2,...,r_n\) can be written in factored form as \[(x-r_1)(x-r_2)...(x-r_n)\]which is why you can do what the other posters suggest. The product of those terms will be 0 at every point where \(x\) is equal to one of the roots, and nowhere else.